OpenStudy (anonymous):

Solve x^2 - 8x + 3 = 0 by completing the square. Which equation is used in the process? a.(x- 4)^2= 0 b.(x- 4)^2= 13 c. (x- 4)^2= 19

OpenStudy (anonymous):

c?

OpenStudy (anonymous):

its not c/?

OpenStudy (shane_b):

Maybe seeing the steps will help: \[x^2-8x+3=0\]\[x^2-8x+4^2+3=16\]\[x^2-8x+4^2=13\]Factor that and you'll get:\[(x-4^2)=13\] Any questions on how that was done?

OpenStudy (anonymous):

nope. got it.

OpenStudy (anonymous):

so its b.

OpenStudy (anonymous):

hey you sure its not b?

OpenStudy (shane_b):

That "Nope" got posted late...Open Study giving me issues. It IS b....it's NOT C.

OpenStudy (shane_b):

The real question is, do you understand the method?

OpenStudy (anonymous):

ooh. ok. yeah i got it.

OpenStudy (anonymous):

thanks.

OpenStudy (shane_b):

Great :)